OpenStudy (anonymous):

Two cards are drawn from a fifty-two-card deck. What is the probability that the draw will yield an ace and a face card?

OpenStudy (mathmale):

Let me turn the question around and ask you a few questions first: (1) How many aces are there in a deck of 52 cards? (2) What is a face card? (3) How many face cards are there, altogether? Your answers, if correct, will help you to solve this problem. Good luck!

OpenStudy (anonymous):

4 aces

OpenStudy (anonymous):

A face card is jack queen and king

OpenStudy (anonymous):

12 face cards altogether

OpenStudy (mathmale):

That's great; thank you. Let's look at the original question: "Two cards are drawn from a fifty-two-card deck. What is the probability that the draw will yield an ace and a face card?" First of all, what's the P that if you draw one card from a 52-card deck, that that card will be an ACE? Secondly, what's the P that if you draw one card from a 52-card deck, that that card will be a FACE card?

OpenStudy (anonymous):

2/52

OpenStudy (mathmale):

BigTime: I'd really like to see your work, not just a possible answer. First of all, what's the P that if you draw one card from a 52-card deck, that that card will be an ACE? Show your work.

OpenStudy (anonymous):

4/52 for your first question

OpenStudy (mathmale):

Good. Secondly, what's the P that if you draw one card from a 52-card deck, that that card will be a FACE card?

OpenStudy (mathmale):

Hint: Bigtim10 A face card is jack queen and king 17 minutes ago 50 Bigtim10 Best Response Medals 0 12 face cards altogether

OpenStudy (mathmale):

Secondly, what's the P that if you draw one card from a 52-card deck, that that card will be a FACE card?

OpenStudy (mathmale):

Please let me know if you're going to take a break or have something else to do. I'll wait for you, but only for a limited time.

OpenStudy (anonymous):

1/52

OpenStudy (anonymous):

Yeah I know what a face card is

OpenStudy (anonymous):

ok

OpenStudy (mathmale):

But you told me there are 12 face cards; how could the probability obtaining a face card now be 1/52?

OpenStudy (anonymous):

My mad I read the question

OpenStudy (mathmale):

So the P that from a deck of 52 cards, you'll pick a FACE card, is what?

OpenStudy (anonymous):

12/52

OpenStudy (mathmale):

Very nice. So: P(ace)m = 4/52 or 1/13, right? P(face card) = 12/52, or 3/13, right?

OpenStudy (anonymous):

yep

OpenStudy (mathmale):

OK! Now we're going to do "sampling with replacement." You'll draw one card, look at it, remember what it is, and then put it back into the deck and shuffle the deck. that way, the P of your obtaining a face card on the second draw is completely independent of whether or not you obtain any particular card on the first draw. Can you agree with that? If not, explain why.

OpenStudy (anonymous):

sure

OpenStudy (mathmale):

Very sorry for the delay. I lost my internet connection. :( So, as you say, the P you'll get an ACE on the first draw is 1/13. If you actually draw an ace and then put the card back and shuffle the deck again, and then draw another card, your chances of obtaining a face card are 4/52. So, P(ace AND face, with replacement, is simply (1/13)(4/52), or (1/13)(1/13)=1/169. Have you seen any thing like this in class, in your online learning materials, or elsewhere?

OpenStudy (anonymous):

no

OpenStudy (mathmale):

What we're calculating here is "joint probability:" It's the probability that on two draws with replacement, the probability P( Ace AND FACE ) = P(Ace) + P(Face). Please go back up, find P(Ace) and P(Face) and calculate P(ACE)*(FACE).

OpenStudy (mathmale):

The following was already part of our conversation, and thus easily available to you: P(ace) = 4/52 or 1/13, right? P(face card) = 12/52, or 3/13, right?

OpenStudy (mathmale):

Please go back up, find P(Ace) and P(Face) and calculate P(ACE)*(FACE). Label your result with " P( ACE AND FACE) = " Let me know just how clear (or unclear) this problem is to you at this point.